# Undergraduate Course Allocation

### v1

@@ D.3 Department-First Priorities
With a department-first priority structure, the Deferred Acceptance mechanism coincides with the Deferred Acceptance mechanism with minority reserves of Hafalir, Yenmez and Yildirim (2013). As in Example D1, this is because students eligible for reserved seats receive the highest priority for all course seats. Hence, we only consider the Deferred Acceptance mechanisms with single and multiple tie-breakings in our simulations. We also drop the analysis of Random Serial Dictatorship with actual course reserves, as its performance is subpar compared to all the other mechanisms. The results of our simulations are presented in Tables D.6D.8.
@@
Table D.6: Each student’s preferred mechanism by year of study under department-first priority structure.
Pseudo-Market with Priorities Deferred Acceptance with single tie-breaking Deferred Acceptance with multiple tie-breakings
[…]
Year 1 14.8 (2.1) 5.5 (1.4) 11.4 (1.5) 1.2 (0.6) 19.4 (1.5) 10.5 (1.1)
Year 2 16.3 (1.7) 9.4 (1.1) 9.3 (1.3) 4.1 (1.2) 20.1 (1.2) 16.3 (1.3)
Year 3 19.5 (2.4) 13.3 (1.3) 8.1 (1.5) 5.6 (1.3) 21.9 (1.4) 20.1 (1.4)
Year 4 14.7 (1.4) 13.2 (1.5) 0.5 (1.0) 6.2 (1.3) 18.4 (1.0) 18.6 (1.3)
A-CEEI A-CEEI with kludgy pricing Probabilistic Serial with Seniority and Reserves
[…]
Year 1 71.1 (1.1) 12.8 (0.7) 62.9 (1.5) 10.0 (0.6) 24.3 (2.3) 25.0 (2.4)
Year 2 73.8 (0.9) 13.8 (0.7) 66.4 (1.3) 11.5 (1.0) 20.6 (2.9) 44.3 (1.6)
Year 3 56.8 (1.2) 24.0 (1.2) 51.4 (1.4) 19.5 (1.0) 33.9 (1.8) 35.9 (1.7)
Year 4 27.9 (1.0) 35.6 (1.3) 26.2 (1.3) 26.7 (1.3) 30.1 (1.1) 29.4 (1.2)
@@
Notes: This table reports the percentage of students who strictly prefer a given mechanism to the benchmark of the Random Serial Dictatorship with optimal course reserves and the percentage of students who strictly prefer the benchmark under the department-first priority structure. Results are averages across $100$ runs with different random component draws.
Table D.6 exhibits each students’ preferred mechanism by year of study under the depart- ment-first priority structure. As opposed to the hybrid priority structure, all students eligible for reserved seats are prioritized above more senior but ineligible students. As a result, a larger percentage of junior students (Years 1 and 2) prefer the PMP, DA-STB, DA-MTB, and A-CEEI with kludgy pricing mechanisms over the RSD benchmark. The comparison is the opposite for senior students (Years $3$ and $4$), with the DA-STB and DA-MTB mechanisms underperforming the RSD benchmark for Year $4$ students. Hence, switching from the RSD benchmark to one of the alternative mechanisms would lead to the redistribution of student welfare from senior to junior students.
Table D.7: The change in the standard deviation of students’ utilities by year of study under a department-first priority structure.
Pseudo-Market with Priorities Deferred Acceptance with single tie-breaking Deferred Acceptance with multiple tie-breaking
Year 1 -0.4% 1.1% -0.02%
Year 2 -0.6% 0.5% -0.1%
Year 3 -2.9% 0.1% -1.7%
Year 4 -2.9% -0.1% -1.5%
A-CEEI A-CEEI with kludgy pricing Probabilistic Serial with seniority and reserves
Year 1 -24.4% -21.4% -4.5%
Year 2 -18.9% -16.2% -7.3%
Year 3 -8.1% -8.6% -9.3%
Year 4 -17.6% -11.7% -9.2%
Notes: The change in the standard deviation of student utility by year of study for the considered mechanisms compared to the Random Serial Dictatorship with optimal course reserves. The outcomes of the Deferred Acceptance mechanism with minority reserves coincides with Deferred Acceptance mechanism both for single and multiple tie-breakings. Results are averages across $100$ runs with different random component draws.
@@
Table D.7 reports each mechanism’s change in the standard deviation of students’ utilities relative to the RSD benchmark. The largest improvements in standard deviation of students’ utilities happen for the A-CEEI, A-CEEI with kludgy pricing, and Probabilistic Serial with seniority and reserves mechanisms. Table D.9 shows that this happens at the expense of priority violations, which more than $50\%$ of students experience in each of these three mechanisms. As in the hybrid priority results, this is a larger percentage of priority violations than we observe in the benchmark ($32.5\%$). In turn, the PMP, DA-STB, and DA-MTB mechanisms have no priority violations, but perform similarly to the RSD benchmark in terms of the standard deviation of students’ utilities.
@@
Table D.8 shows that the percentage of students experiencing schedule envy in each mechanism is similar in comparison to the hybrid priority structure (see Tables 8 and D.4). The PMP, A-CEEI, and A-CEEI with kludgy pricing mechanisms deliver schedule envy bounded by a single course toward students of weakly lower priority, with only $8.2\%$, $4.9\%$, and $4.1\%$ experiencing any envy, respectively. The DA-STB and DA-MTB mechanisms each result in about $13\%$ of students experiencing schedule envy, with about $2\%$ experiencing envy for several courses in the DA-STB mechanism and under $1\%$ in DA-MTB. The RSD with optimal course reserves mechanism has $14.5\%$ of students experiencing schedule envy, with about $2\%$ experiencing envy for several courses. Just $1\%$ of students experience schedule envy in the random assignment of the Probabilistic Serial with seniority and reserves mechanism.
@@
Table D.8: The percentage of students who experience schedule envy under department-first priority structure.
@@
Notes: This table reports the percentage of students who prefer the schedule of a student with weakly lower priority in each course under a department-first priority structure. The first column provides information about students who experience no envy. The other columns show the percentage of students who experience schedule envy bounded by $1,...,5$ courses. Results are averages across $100$ runs with different random component draws.
Table D.9: The percentage of students with a priority violation under a department-first priority structure.
@@
Notes: This table reports the percentage of students whose course schedule can be improved by taking a seat occupied by a student of strictly lower priority under a department-first priority structure. Results are averages across $100$ runs with different random component draws.
@@ D.4 Robustness: Student Choice Sets and Noise Term
This section performs some additional robustness checks. We take the calibrated student utility model calibrated and consider changes in two parameters: the size of student choice sets ($60$, $70$, $80$, $90$), and the standard deviation of the random component of student utility ($\sigma=0.75,1,1.25,1.5$). Figure D.1 reports the performance of the PMP mechanism compared to the RSD benchmark. The graphs show the percentages of students who are indifferent (solid line), prefer the PMP mechanism (dashed line), and prefer the RSD benchmark (dotted line). In contrast to the main text, the reported simulation results are based on $50$ random component draws.
The top four graphs show comparative statics results for the size of student choice sets. The percentage of students indifferent between the PMP and RSD mechanisms is relatively constant across all sizes ($74.8\%$ to $76.3\%$), slightly lower for student choice sets of size $90$ ($68.5\%$). A relatively equal percentage of Year $1$ students strictly prefer each mechanism. For Year $2$ through Year $4$ students, a larger percentage of students prefer the PMP mechanism regardless of the size of the choice set.
The bottom four graphs show comparative statics results for the standard deviation of the noisy component $\sigma$. The percentage of students indifferent between the PMP mechanism and RSD mechanism significantly declines from $\sigma=0.75$ to $\sigma=1.5$, with the largest decrease for Year 1 students ($85.7\%$ to $13.8\%$). Again, a relatively equal percentage of Year $1$ students strictly prefer each mechanism. For Year $2$ through Year $4$ students, a larger percentage of students prefer the PMP mechanism regardless of the size of the standard deviation. The benefits of the PMP mechanism are also larger for more noisy environments, where student preferences for the same schedules could be quite different.
@@ E.1 Program
Most simulations have been done using Mathematica 14.0.0.0 on a server with 12 Cores of CPU (2.6 GHz of an Intel Xeon Gold 6126 CPU), 128 GB of RAM, 128 GB of SSD hard drive space, and a Windows Server 2019 Standard operating system. There are three simulation programs:
@@
    • RandomParam generates the random student choice sets and random utility components.
@@
The programs simulations.nb and simulations-appendix.nb have the same sections and employ similar functions. We label the functions only used in simulations-appendix.nb with an asterisk.
@@
  • Preferred Mechanism Results produces results for Tables 5, D.2, and D.6.
  • Envy Results produces results for Tables 8, D.4, and D.8.
  • Priority Violations produces results for Tables D.5 and D.9.
  • Utility Results loads in pricing data from pmp-data.txt, aceei-data.txt*, and kludgy-data.txt* and produces results for Tables 6, 7, D.1, D.3, and D.7.
@@ E.2 Computational Algorithm
We developed an algorithm to find a Pseudo-Market Equilibrium with Priorities, taking in student utilities $v$, student priorities $priorities$, and a tie-breaking order over students $studentrank$ and outputting an allocation $bestdemand$ and a vector of course prices $bestt$. Its main stages are described below, with the pseudocode shown on page Algorithm.
@@
Initialization. We first allocate budgets of fake money to students. To do so, we reorder the set of student budgets $b$, evenly spread values from $1$ to $1.251+\beta=$ 1.25, according to $studentrank$. The function PMPInitialPrice takes in student utilities $v$ and provides an initial estimate for the price vector based on students’ ideal bundles of courses initdemand. [^40] A student’s ideal bundle is the set of courses they would demand if all course prices were zero. Rather than working in $M\!R$-dimensional space, we leverage the approach in the proof of Theorem 1 and parameterize priority-specific prices as $t\in[0,R\overline{b}]^{M}$ We collect the main parameters of the algorithm in $pmpparam\leftarrow\{v,b,priorities,initdemand\}$ and initialize the variables $boundcheck\leftarrow 0$ and $oversubcheck\leftarrow 0$. These last two parameters serve as exit flags for Phases I and II of the algorithm described below.
@@
The main innovation of our search process is that course demand is only updated for two subsets of students at each step. The first subset includes students who can no longer afford the allocation determined in the previous step. The second includes students who did not demand their ideal bundle in the previous step and have a strictly positive utility for courses with recently decreased prices, where these prices are within the student’s budget. For any student outside these two subsets, their optimal course demand remains unchanged between consecutive steps of the search process. This innovation has proven particularly effective for large student populations, significantly reducing the time required for the search process compared to having all students update their demand at each step.
Phase II. The second phase starts with the best price of Phase I and works to reduce the excess course demand to levels below those observed in the data. For each course with positive excess demand and each student who demands it, we use a dual program to calculate the minimum price change needed so that the student no longer demands the course. We then sort these price changes and apply the smallest price change that fully eliminates the excess demand for the course. For example, if a course’s excess demand is 2, we apply the second smallest price change among students who demand the course. We repeat this process until two conditions are met: (i) the distribution of excess demand errors stochastically dominates the distribution observed in the data, and (ii) no course is oversubscribed by more than $k-1=4$ seats. If both (i) and (ii) are satisfied we set $oversubcheck=1$. We also update the value of $boundcheck$ as the market-clearing error changes. The output of Phase II includes the price vector $bestt$ and the exit flags $boundcheck$ and $oversubcheck$.
Final Allocation. After Phase II, we provide checks that the optimization problem for each student at the final prices is correctly solved and produce the final course allocation. If errors occur, the FinalAllocation function will correct them. If errors result in a violation of Phase I and II conditions, we return $bestdemand=\emptyset$. If such an event happens, we reset the exit flags and the number of rounds to the maximum value and restart the process with a new initial price vector. The last step is done using the last IF statement.
Last, we describe the two while loops in the algorithm. The outer while loop continues as long as $boundcheck=0$ or $oversubcheck=0$, allowing the search for a price vector to proceed. At each step, we add a small amount of noise to the initial price vector to begin the search from a slightly different starting point, and we reset the number of rounds in the inner loop to zero. The inner loop has the same exit conditions, with an additional condition on the number of rounds. If either $boundcheck=0$ or $oversubcheck=0$ after Phase II, we repeat the process. We allow up to $maxrounds=6$ iterations of Phase I and II within the inner loop making the search process more persistent during additional rounds. If the algorithm fails to find a solution after $maxrounds$ iterations, we restart the process with a new initial price vector. Simulations indicate that typically only 1-4 iterations occur between Phases I and II in the inner loop. We did not encounter cases requiring a restart through the outer loop for the main simulations. The simulations in the appendix Section D.4 checking the robustness of our results do encounter some challenging cases where a restart is necessary.
Algorithm Pseudo-Market with Priorities
1:student utilities $v$, student priorities $priorities$, student ranking $studentrank$
2:course allocation $bestdemand$, prices $bestt$
4:$pmpparam\leftarrow\{v,b,priorities,initdemand\}$, $boundcheck\leftarrow 0$, $oversubcheck\leftarrow 0$

### v2

@@ D.3 Department-First Priorities
Figure D.4: Each student’s preferred mechanism under the department-first priority structure.
@@
Notes: This figure reports the percent of students who strictly prefer each given mechanism to the Random Serial Dictatorship with optimal course reserves benchmark and the percent of students who strictly prefer the benchmark under the department-first priority structure. Results are averages across 100 runs with different random component draws.
With a department-first priority structure, the Deferred Acceptance mechanism coincides with the Deferred Acceptance mechanism with minority reserves of Hafalir, Yenmez and Yildirim (2013). As in Example D1, this is because students eligible for reserved seats receive the highest priority for all course seats. Hence, we only consider the Deferred Acceptance mechanisms with single and multiple tie-breakings in our simulations. We also drop the analysis of Random Serial Dictatorship with actual course reserves, as its performance is subpar compared to all the other mechanisms. The results of our simulations are presented in Figures D.4D.6 and Tables D.5D.6.
Figure D.5: Improvements in utility among students with changing schedules under the department-first priority structure.
Notes: This figure reports the percent improvement in mean utility among students with changing schedules for each given mechanism over the Random Serial Dictatorship with optimal course reserves benchmark under the department-first priority structure. Results are averages across $100$ runs with different random component draws.
Figure D.4 displays the results on each student’s preferred mechanism under the department-first priority structure. We can observe that an aggregate relative comparison is similar to the results for the hybrid priority structure (Figures 1 and D.1). In particular, there is a larger percentage of students that strictly prefer each of the given mechanisms than students who prefer the RSD benchmark. The only exception is the PS mechanism, which is consistent across both priority structures.
As opposed to the hybrid priority structure, the department-first priority structure has all students eligible for reserved seats prioritized above more senior but ineligible students. For PMP, DA-STB, and DA-MTB, this preferential treatment benefits more junior students (Years $1$ and $2$) and harms senior students (Years $3$ and $4$). This is illustrated in Figure D.5 by the magnitudes in improvements in utility among students with changing schedules. [^38] On average, the PMP mechanism changes schedules for 317, 414, 466, and 397 students in Years 1, 2, 3, and 4, respectively; the DA-STB mechanism does so for 198, 216, 196, and 169 students; the DA-MTB mechanism does so for 468, 586, 599, and 527 students; the A-CEEI mechanism does so for 1313, 1412, 1149, and 904 students; the A-CEEI with kludgy pricing mechanism does so for 1140, 1254, 1008, and 753 students; and the PS mechanism does so for 772, 1044, 992, and 848 students. The results are less sharp for the A-CEEI and A-CEEI with kludgy pricing mechanisms. These do not respect course priorities, leading to the benefit being spread across years of study. We also observe the opposite effect for the PS mechanism: an improvement in utility is larger for more senior students compared to those who are more junior.
@@
Figure D.6 reports each mechanism’s change in the standard deviation of students’ utilities relative to the RSD benchmark. The largest improvements happen for the A-CEEI, A-CEEI with kludgy pricing, and Probabilistic Serial with seniority and reserves mechanisms. Table D.6 shows that this happens at the expense of priority violations, which more than $50\%$ of students experience in each of these three mechanisms. As in the hybrid priority structure results, this is a larger percentage of priority violations than we observe in the benchmark ($32.5\%$). In turn, the PMP, DA-STB, and DA-MTB mechanisms have no priority violations, but perform closer to the RSD benchmark in terms of the standard deviation of students’ utilities.
Figure D.6: The change in the standard deviation of students’ utilities under the department-first priority structure.
Notes: This figure reports the percent change in the standard deviation of students’ utilities compared to the Random Serial Dictatorship with optimal course reserves benchmark under the department-first priority structure. Results are averages across 100 runs with different random component draws.
@@
Table D.5 shows that the percentage of students experiencing schedule envy in each mechanism is similar in comparison to the hybrid priority structure (see Tables 5 and D.3). The PMP, A-CEEI, and A-CEEI with kludgy pricing mechanisms deliver schedule envy bounded by a single course toward students of weakly lower priority, with only $8.2\%$, $4.9\%$, and $4.1\%$ experiencing any envy, respectively. The DA-STB and DA-MTB mechanisms each result in about $13\%$ of students experiencing schedule envy, with about $2\%$ experiencing envy for several courses in the DA-STB mechanism and under $1\%$ in DA-MTB. The RSD with optimal course reserves mechanism has $14.5\%$ of students experiencing schedule envy, with about $2\%$ experiencing envy for several courses. Just $1\%$ of students experience schedule envy in the random assignment of the Probabilistic Serial with seniority and reserves mechanism.
@@
Table D.5: The percentage of students who experience schedule envy under the department-first priority structure.
@@
Notes: This table reports the percentage of students who prefer the schedule of a student with weakly lower priority in each course in each mechanism under the department-first priority structure. The first column includes students who experience no envy. The other columns display students who experience schedule envy bounded by 1,,51,...,5 courses. Results are averages across $100$ runs with different random component draws.
Table D.6: The percentage of students with a priority violation under the department-first priority structure.
@@
Notes: This table reports the percentage of students whose course schedule can be improved by taking a seat occupied by a student of strictly lower priority for each mechanism under the department-first priority structure. Results are averages across $100$ runs with different random component draws.
@@ D.4 Robustness: Student Choice Sets and Noise Term
This section performs some additional robustness checks. We take the calibrated student utility model and consider changes in two parameters: the size of student choice sets ($60$, $70$, $80$, $90$) and the standard deviation of the random component of student utility ($\sigma=0.75,1,1.25,1.5$). Figure D.7 reports the performance of the PMP mechanism compared to the RSD benchmark. The graphs show the percentages of students who are indifferent (solid line), prefer the PMP mechanism (dashed line), and prefer the RSD benchmark (dotted line). Reported simulation results are based on $50$ random component draws.
The top four graphs show comparative statics results for the size of student choice sets. The percentage of students indifferent between the PMP and RSD mechanisms is relatively constant across all sizes ($68.5\%$ to $76.3\%$), slightly lower for student choice sets of size $90$. A relatively equal percentage of Year $1$ students strictly prefer each mechanism. For Year $2$ through Year $4$ students, a larger percentage of students prefer the PMP mechanism regardless of the size of the choice set.
The bottom four graphs show comparative statics results for $\sigma$. The percentage of students indifferent between the PMP mechanism and RSD mechanism significantly declines from $\sigma=0.75$ to $\sigma=1.5$, with the largest decrease for Year 1 students ($85.7\%$ to $13.8\%$). Again, a relatively equal percentage of Year $1$ students strictly prefer each mechanism. For Year $2$ through Year $4$ students, a larger percentage of students prefer the PMP mechanism regardless of the size of the standard deviation. The benefits of the PMP mechanism are larger for noisier environments.
@@ E.1 Program
Most simulations have been done using Mathematica 14.0.0.0 on a server with 12 Cores of CPU (2.6 GHz of an Intel Xeon Gold 6126 CPU), 128 GB of RAM, 128 GB of SSD hard drive space, and a Windows Server 2019 Standard operating system. There are three programs:
@@
    • RandomParam generates random student choice sets and utility components.
@@
Programs simulations.nb and simulations-appendix.nb have the same sections and similar functions. We label functions only used in simulations-appendix.nb with an asterisk.
@@
  • Preferred Mechanism Results produces results for Figures 1, D.1, and D.4.
  • Envy Results produces results for Tables 5, D.3, and D.5.
  • Priority Violations produces results for Tables D.4 and D.6.
  • Utility Results loads in data from pmp-data.txt, aceei-data.txt*, and kludgy-data.txt* and produces results for Tables D.1 and D.2 and Figures 2, 3, D.2, D.3, D.5, and D.6.
@@ E.2 Computational Algorithm
We developed an algorithm to find a Pseudo-Market Equilibrium with Priorities, taking in student utilities $v$, student priorities prioritiespriorities, and a tie-breaking order over students studentrankstudentrank, and outputting an allocation bestdemandbestdemand and a vector of course prices besttbestt. Its main stages are described below, with the pseudocode shown on page Algorithm.
@@
Initialization. We first reorder the set of student budgets $b$, evenly spread values from $1$ to $1+\beta=1.25$, and assign them to students in order of studentrankstudentrank. The function PMPInitialPrice takes in student utilities $v$ and provides an initial estimate for the price vector based on students’ most-preferred course schedules initdemand. Rather than working in MRM\!R-dimensional space, we leverage the approach in the proof of Theorem 1 and parameterize priority-specific prices as $t\in[0,R\overline{b}]^{M}$ . We collect these parameters in $pmpparam\leftarrow\{v,b,priorities,initdemand\}$ and initialize $boundcheck\leftarrow 0$ and $oversubcheck\leftarrow 0$, which serve as exit flags for Phases I and II of the algorithm.
Phase I. The first phase of the algorithm is a $\hat{\text{a}}t$onnement process, searching for a price vector by repeatedly adjusting prices proportionally to excess demand. If the market-clearing error remains above the theoretical bound THEORYBOUND $\approx 43.5$, this phase ends when there is no improvement in the error over a fixed number of consecutive steps. If the market-clearing error is below the theoretical bound, we stop once there is no substantial improvement (more than $1\%$) in error over the same number of consecutive steps. The algorithm sets the flag $boundcheck=1$ if the error falls below THEORYBOUND and $boundcheck=0$ otherwise. As output, Phase I returns the price vector besttbestt and the exit flag boundcheckboundcheck. The upper limit on the consecutive steps with no improvement and some other parameters of Phase I depend on nroundnround making the search process more persistent for later rounds.
The main innovation in Phase I is that demands are only updated for two subsets of students in each step. The first includes students who can no longer afford the allocation determined in the previous step. The second includes students who did not demand their most-preferred course schedule in the previous step and have a positive utility for courses with recently decreased prices (within the student’s budget). Any student outside these two subsets has her demand unchanged between steps. This has proven particularly effective, significantly shortening the search process compared to updating every student’s demand.
Phase II. The second phase starts with the best price of Phase I and reduces excess course demand to levels below those observed in the data. For each positive excess demand course and each demanding student, we use a dual program to calculate the minimum price change needed so the student no longer demands the course and apply the smallest change that fully eliminates the course’s excess demand. We repeat this process until two conditions are met: (i) the distribution of excess demand errors stochastically dominates the distribution observed in the data, and (ii) no course is oversubscribed by more than k1=4k-1=4 seats. If both (i) and (ii) are satisfied we set $oversubcheck=1$. We also update the value of boundcheckboundcheck as the market-clearing error changes. The output of Phase II includes price vector besttbestt and the exit flags boundcheckboundcheck and oversubcheckoversubcheck.
Final Allocation. After Phase II, we check each student’s demand is correctly computed at the final prices and produce the final course allocation. If errors occur, the FinalAllocation function will correct them. If errors violate the Phase I and II conditions, we set $bestdemand=\emptyset$ and restart the process with a new initial price vector.
The outer while loop of the algorithm continues as long as $boundcheck=0$ or $oversubcheck=0$. At each step, we add a small amount of noise to the initial price vector to begin the search from a slightly different starting point. The inner loop has the same exit conditions, with an additional condition on the number of rounds. If either boundcheckboundcheck or $oversubcheck=0$ after Phase II, we repeat the process. We allow up to $maxrounds=6$ iterations of Phase I and II within the inner loop making the search process more persistent during additional rounds. If the algorithm fails to find a solution after maxroundsmaxrounds iterations, we restart with a new initial price vector. Typically only 1-4 iterations occur between Phases I and II in the inner loop. We did not encounter cases requiring a restart through the outer loop for the main simulations. The simulations in the appendix Section D.4 checking the robustness of our results do encounter some challenging cases where a restart is necessary.
Algorithm Pseudo-Market with Priorities
1:student utilities $v$, student priorities prioritiespriorities, student ranking studentrankstudentrank
2:course allocation bestdemandbestdemand, prices besttbestt